40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Mar 19 15:34:41 2020, anchal, DailyProgress, North Cavity, North path's buggy nature solved IMG_20200319_154013.jpg
    Reply  Sun Mar 22 17:23:29 2020, anchal, DailyProgress, North Cavity, North path's buggy nature NOT solved 
       Reply  Wed Mar 25 15:50:57 2020, anchal, DailyProgress, North Cavity, North path's buggy nature NOT solved Screen_Shot_2020-03-25_at_3.48.24_PM.png
          Reply  Tue Apr 14 12:24:16 2020, anchal, DailyProgress, North Cavity, North path's buggy nature Strikes Again! NorthPathWentBuggy.pdf
             Reply  Thu Apr 16 18:08:56 2020, anchal, DailyProgress, North Cavity, North path stopped being buggy NorthPathStoppedBeingBuggy.pdf
Message ID: 2565     Entry time: Wed Mar 25 15:50:57 2020     In reply to: 2564     Reply to this: 2567
Author: anchal 
Type: DailyProgress 
Category: North Cavity 
Subject: North path's buggy nature NOT solved 

Today, magically almost, the North Path was found to be locking nicely without the noise. I was waiting for the beatnote to reach the detector's peak frequency when in about 40 min, it started going haywire again. No controls were changed to trigger any of this and as far as I know, nobody entered the lab. Something is flakey or suddenly some new environmental noise is getting coupled to the experiment. Attached is the striptool screenshot of the incident and the data dumped. In the attached screenshot, channel names are self-explanatory, all units on y axis are mW on left plot (note the shifted region, but same scale of North Cavity Transmission Power) and MHz on the right plot for Beatnote Frequency.

I know for sure that everything until PMC is good since when only PMC is locked, I do not see huge low frequency noise in the laser power transmitted or reflected from the PMC. But whatver is this effect, it makes the FSS loop unstable and eventually it unlocks, then locks again and repeats.


Evidences

Attachment 1: Screen_Shot_2020-03-25_at_3.48.24_PM.png  79 kB  Uploaded Wed Mar 25 17:15:48 2020  | Hide | Hide all
Screen_Shot_2020-03-25_at_3.48.24_PM.png
ELOG V3.1.3-